OpenStudy (anonymous):

3/4 / -3/7 = = -1/28 right ?

OpenStudy (anonymous):

no ;)

OpenStudy (tkhunny):

\(\dfrac{\dfrac{3}{4}}{\dfrac{-3}{7}} = \dfrac{3}{4}\cdot\dfrac{7}{-3}\) Rethink your response?

OpenStudy (nurali):

\[\frac{ 3 }{ 4 }*-\frac{ 7 }{ 3 }\]

OpenStudy (mathstudent55):

\( \dfrac{\dfrac{3}{4}}{\dfrac{-3}{7}} \) \( = \dfrac{3}{4} \div \left(-\dfrac{3}{7} \right)\) \( = \dfrac{3}{4} \times \left(-\dfrac{7}{3} \right)\) \( = -\dfrac{7}{4} \)